OpenStudy (anonymous):

Having problems with multi step equations, solving for x, or in this case b, with two sided equations. How to do this? 8 - 2 x b = -2/3 x (12 x b + 6)

OpenStudy (texaschic101):

8 - 2b = -2/3(12b + 6) --- Is this your problem ??

OpenStudy (anonymous):

Yes

OpenStudy (texaschic101):

first we start on the right side and distribute the -2/3 through the parenthesis. Can you do that ?

OpenStudy (anonymous):

By distribute, you mean multiply?

OpenStudy (anonymous):

8 - 2b = -2/3(12b + 6 that's how it goes

OpenStudy (texaschic101):

yes...multiply everything in the parenthesis by -2/3

OpenStudy (anonymous):

How would you do this without using the variable?

OpenStudy (netlopes1):

\[8-2b=-\frac{2}{3}(12b+6) \rightarrow\]

OpenStudy (netlopes1):

is this?

OpenStudy (anonymous):

I mean leaving it out, it would be -12.

OpenStudy (anonymous):

Yes @netlopes1

OpenStudy (texaschic101):

8 - 2b = (-2/3) * 12b + (-2/3) * 6 8 - 2b = -24/3b - 12/3 ---- reduce when you can 8 - 2b = -8b - 4 are you still following this ??

OpenStudy (anonymous):

Not completely, you replaced b with -2/3 :P

OpenStudy (netlopes1):

\[8-2b=-\frac{2}{3}.12b - \frac{2}{3}.6 \rightarrow\]

OpenStudy (texaschic101):

where you see that ?? I did not replace b with -2/3

OpenStudy (netlopes1):

This is a distributive rule, ok?

OpenStudy (anonymous):

Ohh, you left b but added -2/3.

OpenStudy (netlopes1):

\[8-2b=-\frac{24b}{3}-\frac{12}{3}\rightarrow do\quad you\quad understand\quad?\]

OpenStudy (texaschic101):

to get rid of the parenthesis we use the distributive property...that means you take the number right before the parenthesis and multiply it by everything in the parenthesis. 8 - 2b = -2/3 (12b + 6) = 8 - 2b = (-2/3 * 12b) + (-2/3 * 6) = 8 - 2b = (-24/3)b + (-12/3) = 8 - 2b = -8b - 4 any better understand this ??

OpenStudy (netlopes1):

ok, texaschic101

OpenStudy (netlopes1):

\[finally,\quad 8-2b=-8b-4\rightarrow6b=-12\rightarrow b=-2\]

OpenStudy (texaschic101):

you understand ?

OpenStudy (netlopes1):

\[6-3b=-\frac{3}{2}(-4b+4) \rightarrow\]

OpenStudy (netlopes1):

If you understand, answer this equation, ok? It's the same "model", I mean.